OpenStudy (anonymous):

what is the fraction of 2.4?

OpenStudy (anonymous):

2 and 2/5

OpenStudy (anonymous):

nope its not possible. its not an answer on my worksheet. i just need help finding the answer. please try again.

OpenStudy (anonymous):

well maybe it is in improper form, in which case it would be 12/5

OpenStudy (anonymous):

im sorry but no thats not the case.

OpenStudy (anonymous):

the percent is 2/4% and i have 2 find the decimal and fraction

OpenStudy (anonymous):

Ah, I thought it was just 2.4, not 2.4%. 2.4% is just 2.4/100, which is 24/1000=12/500=6/250=3/125